OpenStudy (anonymous):

Use inductive reasoning to describe the pattern. Then find the next two numbers in the pattern. –3, 9, –27, 81, . . . subtract 12 from the previous term; –243, 729 multiply the previous term by 3; 243, 729 multiply the previous term by –3; –243, 729 add 12 to the previous term; 69, 57

OpenStudy (anonymous):

-3 to 9 = multiply by -3 9 to -27 = multiply by -3 -27 to 81 = multiply by -3 next to numbers are: 81 x -3 = -243 -243 x -3 = 729

OpenStudy (anonymous):

Choice A. Does -3 - 12 = 9? Does 9 - 12 = -27? Choice B. Does -3 x 3 = 9 and does 9 x 3 = -27? Choice C. Does -3 x -3 = 9 and does 9 x 3 = -27? Choice D. Does -3 + 12 = 9 and does 9 + 12 =-27??